OpenStudy (anonymous):

The excursion boat on the river takes 2½ hours to make the trip to a point 12 miles upstream and to return. If the rate at which the boat travels in still water is 5 times the rate of the river current, what is the rate of the current? Which of the following equations can be used to solve for c, the rate of the current?

OpenStudy (anonymous):

Let the current speed be "c"; Then the boat speed in still wate is "5c": ----------------------- Upstream speed: distance = 12 miles ; rate = 5c-c = 4c mph ; time = 12/4c = 3/c hrs --------------------------- Downstream speed: distance = 12 miles ; rate = 5c+c = 6c mph ; time = 12/6c = 2/c hrs --------------------------- Equation: time + time = 2 1/2 hrs 3/c + 2/c = 5/2 Multiply thru by 2c to get: 6 + 4 = 5c c = 2 mph (speed of the current) 5c = 10 mph (speed of the boat in still water) ==================================================